Find the Laplace transforms of the following function:

10. L{e^-t}


1
Expert's answer
2022-05-24T08:27:04-0400
"F(s)=L(e^{-t})=\displaystyle\int_{0}^{\infin}e^{-st}e^{-t}dt"

"=\lim\limits_{A\to\infin}\displaystyle\int_{0}^{A}e^{-(s+1)t}dt=[-\dfrac{e^{-(s+1)t}}{s+1}]\begin{matrix}
 A \\
 0
\end{matrix}"

"=-0+\dfrac{1}{s+1}=\dfrac{1}{s+1}"

"L(e^{-t})=\dfrac{1}{s+1}"